### v1

# CRMAgent: A Multi-Agent LLM System for E-Commerce CRM Message Template Generation

Abstract
In e-commerce private-domain channels such as instant messaging and e-mail, merchants engage customers directly as part of their Customer Relationship Management (CRM) programmes to drive retention and conversion. While a few top performers excel at crafting outbound messages, most merchants struggle to write persuasive copy because they lack both expertise and scalable tools. We introduce CRMAgent, a multi-agent system built on large language models (LLMs) that generates high-quality message templates and actionable writing guidance through three complementary modes. First, group-based learning enables the agent to learn from a merchant’s own top-performing messages within the same audience segment and rewrite low-performing ones. Second, retrieval-and-adaptation fetches templates that share the same audience segment and exhibit high similarity in voucher type and product category, learns their successful patterns, and adapts them to the current campaign. Third, a rule-based fallback provides a lightweight zero-shot rewrite when no suitable references are available. Extensive experiments show that CRMAgent consistently outperforms merchants’ original templates, delivering significant gains in both audience-match and marketing-effectiveness metrics.
CRMAgent: A Multi-Agent LLM System for E-Commerce
CRM Message Template Generation
Yinzhu Quan Georgia Institute of Technology Atlanta, GA 30332, USA yquan9@gatech.edu Xinrui Li ByteDance Inc. Seattle, WA 98004, USA ava.li@bytedance.com Ying Chen ByteDance Inc. San Jose, CA 95110, USA eric.ychen@bytedance.com
1 Introduction
Customer Relationship Management (CRM) Kumar and Reinartz (2018) messages serve as a primary lever for user retention and revenue generation on modern e-commerce platforms. Merchants routinely push promotional notifications that announce new products, highlight price drops, or deliver loyalty rewards to segmented audiences through app and messaging channels. Crafting effective templates remains a bottleneck because most merchants lack dedicated marketing expertise, work under tight time constraints, and therefore rely on generic system-generated copy. Many CRM messages consequently underperform, leaving substantial engagement and conversion potential untapped.
Recent work has shown the effectiveness of LLM agents for structured decision-making in operational domains such as logistics Ieva et al. (2025); Quan et al. (2025), supply chain management Li et al. (2023); Quan and Liu (2024b); Wang et al. (2025), and economic reasoning Quan and Liu (2024a); Guo and Yang (2024); Liu and Quan (2025a). To address this performance gap in CRM, we present CRMAgent (Figure 1), a multi-agent system built on large language models (LLMs) that automatically improves underperforming CRM message templates in e-commerce. Using historical engagement data, CRMAgent identifies low-performing messages, retrieves high-quality examples from similar contexts, and produces improved variants tailored to each merchant and audience segment group. The workflow is divided among four specialized agents: ContentAgent, RetrievalAgent, TemplateAgent, and EvaluateAgent, each responsible for a distinct subtask.
Our contributions are threefold: (1) We present an end-to-end multi-agent LLM framework for e-commerce CRM message template generation. The system decomposes the task into four specialized agents: ContentAgent, RetrievalAgent, TemplateAgent and EvaluateAgent, which together handle data ingestion, analysis, template generation and automatic quality assessment. (2) To cover diverse data scenarios, we develop a TemplateAgent which can learn from a merchant’s own high-performing messages within the same audience segment (group-based learning), retrieve and adapt cross-merchant exemplars that match audience segment, voucher type and product category (retrieval-and-adaptation), and fall back to a rule-based zero-shot rewrite when no suitable references exist. (3) And we implement EvaluateAgent which supports two evaluation modes: assigning separate scores for audience fit and marketing effectiveness to each candidate, and directly selecting the superior message.
2 Related Work
Recent multi-agent LLM systems adopt modular architectures to address structured generation tasks across diverse domains. For instance, Singh et al. (2024) design role-based agents for collaborative travel planning, while DocAgent Yang et al. (2025) structures code documentation into sequential reader–writer–verifier stages. MAxPrototyper Yuan et al. (2024) coordinates theme, text, and visual elements under a central controller for UI prototyping, and InvAgent Quan and Liu (2024b) demonstrates that agent specialization benefits inventory management in logistics. Similarly, CRMAgent applies a task-specialized multi-agent framework to CRM message rewriting, with agents dedicated to content analysis, retrieval, generation, and evaluation. Complementary to modularity, retrieval-augmented generation (RAG) improves LLM output by grounding in exemplar data or structured knowledge. Retrieve-and-Edit Hashimoto et al. (2018), PEARL Malik et al. (2024), and Chen et al. (2024) leverage retrieval to enhance preference modeling and factual accuracy. Recent work also explores temporal event modeling and retrieval Liu and Quan (2024, 2025b), combining LLMs with temporal point processes to unify event semantics and timing, while Liu et al. (2024) show that fine-tuning with retrieved exemplars improves personalization. Extending these directions, CRMAgent retrieves high-performing CRM templates based on audience, product, and voucher metadata, and generates tailored messages using marketing feedback.
3 Multi-Agent Collaboration for CRM Message Template Generation
This section presents the architecture of the CRMAgent system, detailing the roles and interactions of its four specialized agents.
3.1 ContentAgent
ContentAgent performs structured diagnosis of CRM message templates by comparing high- and low-performing examples within the same audience segment. It summarizes why certain messages achieved better engagement than others, focusing on structural and stylistic patterns in content design. Specifically, it analyzes differences in phrasing, organization, targeting logic, and the way incentives are presented. To ensure contextual relevance, engagement percentiles are calculated within each audience segment, defined by the plan’s audience label. Figure 2 shows the prompt used to guide this comparative analysis.
3.2 RetrievalAgent
RetrievalAgent retrieves high-quality CRM message templates from other merchants within the same audience group, selecting those that are most similar in product category and voucher type. When no suitable high-performing templates exist within the same merchant-audience group, RetrievalAgent searches cross-merchant examples to support rewriting. Each candidate message is represented by a dense embedding generated using a SentenceTransformer (Reimers and Gurevych, 2019) model all-MiniLM-L6-v2 Wang et al. (2020). Embeddings are computed based on structured metadata, including the product category hierarchy and voucher attributes. These vectors are indexed using Faiss Douze et al. (2024) for efficient similarity search. At inference time, the agent encodes the target message’s metadata into a dense vector and retrieves the top- most similar templates to serve as few-shot references for TemplateAgent. If both product and voucher metadata are missing, the system skips RetrievalAgent and falls back to rule-based rewriting.
3.3 TemplateAgent
TemplateAgent rewrites low-performing CRM message templates by combining diagnostic insights from ContentAgent with high-performing exemplars. These exemplars are either sourced from the same merchant and audience segment or retrieved across merchants by RetrievalAgent. The agent receives four types of input: the original message, a structured explanation of its weaknesses, a summary of success patterns, and one or more reference templates. To improve clarity, persuasiveness, and alignment with audience expectations while still preserving the original promotional intent such as product and voucher offers, TemplateAgent follows a three-stage template generation strategy: it first uses in-group exemplars if available, otherwise retrieves top-ranked examples from the same audience group with similar product and voucher characteristics, and finally generates a message using rule-based rewriting when no suitable references are found. The exemplar-guided prompt is shown in Figure 3, and the rule-based fallback prompt is shown in Figure 4.
3.4 EvaluateAgent
EvaluateAgent assesses the quality of CRM message template revisions through two complementary evaluation modes. In the scoring mode, it rates each message independently on two criteria: (1) audience match and (2) marketing effectiveness. Each criterion uses a three-level scale with values of 1 (poor), 3 (average), and 5 (excellent). Concise justifications are provided to support each score and enhance interpretability. In the comparison mode, the agent receives two anonymized versions labeled as A and B and selects the more persuasive and audience-appropriate candidate. The agent is not told which version is original and which is rewritten, ensuring unbiased evaluation based solely on message quality. The prompt for the scoring mode appears in Figure 5, and the prompt for the comparison mode is shown in Figure 6.
4 Experiments
We conduct comprehensive experiments to assess the effectiveness of our proposed multi-agent framework for CRM message template generation. This section is organized into three parts: we first present quantitative evaluation results across audience segments, analyzing audience alignment, marketing quality, and preference metrics. We then perform an ablation study to examine the impact of different LLM configurations for each agent component. Finally, we include qualitative case studies to highlight how generated messages differ from the originals in structure, tone, and targeting strategy.
4.1 Experimental Results
| Audience Segment | Count | Audience Score | Market Score | Preference | |||||
| Ori | Gen | (%) | Ori | Gen | (%) | Preferred | Rate (%) | ||
| Potential New Customers | 701 | 3.32 | 4.71 | 41.87 | 3.14 | 4.78 | 52.23 | Gen | 88.16 |
| New Buyers | 681 | 3.63 | 4.38 | 20.66 | 3.32 | 4.57 | 37.65 | Gen | 80.62 |
| Lapsing Buyers | 515 | 3.84 | 4.12 | 7.29 | 3.41 | 4.38 | 28.45 | Gen | 67.18 |
| Abandoned Cart Buyers | 493 | 4.76 | 4.91 | 3.15 | 3.17 | 4.83 | 52.37 | Gen | 87.22 |
| Unpaid Order Buyers | 439 | 4.95 | 4.89 | -1.21 | 3.30 | 4.73 | 43.33 | Gen | 83.60 |
| Post-Purchase Group | 352 | 4.91 | 4.74 | -3.46 | 3.18 | 4.56 | 43.40 | Gen | 76.70 |
| Price-Drop Group | 267 | 4.99 | 4.79 | -4.01 | 4.03 | 4.39 | 8.93 | Gen | 63.67 |
| Active Old Followers | 186 | 4.06 | 3.89 | -4.19 | 3.25 | 4.60 | 41.54 | Gen | 70.43 |
| Frequent Buyers | 140 | 4.20 | 4.20 | 0.00 | 3.54 | 4.44 | 25.42 | Gen | 64.29 |
| New Followers | 125 | 4.70 | 4.46 | -5.11 | 3.77 | 4.52 | 19.89 | Gen | 79.20 |
| Repeat Buyers | 58 | 4.17 | 4.00 | -4.08 | 3.41 | 4.34 | 27.27 | Gen | 56.90 |
| Overall | 3957 | 4.18 | 4.56 | 9.09 | 3.33 | 4.61 | 38.44 | Gen | 78.44 |
We evaluate the quality of CRMAgent generated templates against merchant originals across 11 audience segments, including potential new customers, new buyers, lapsing buyers, abandoned cart buyers, unpaid order buyers, post-purchase users, price-drop seekers, active old followers, frequent buyers, new followers, and repeat buyers. First, we quantify gains in audience alignment, marketing effectiveness, and overall preference using the structured scores produced by EvaluateAgent. Next, we compute BERTScore-F1 and chrF to verify that the rewritten copy preserves the core semantic content while allowing phrasing changes. Together, these two sets of metrics offer a comprehensive view of both persuasive impact and faithfulness to the original intent.
We use GPT-4o (OpenAI, 2024) as the underlying model for ContentAgent, TemplateAgent, and EvaluateAgent in this experiment. Table 1 shows that the generated templates consistently outperform the originals across most audience segments. Overall, the audience fit score rises by 9.09%, and the marketing effectiveness score jumps by 38.44%. EvaluateAgent chooses the rewritten version in 78.44% of blind comparisons, confirming its superiority. Qualitatively, segments that typically need a stronger push, such as potential new customers, benefit the most: the new copy adds explicit urgency and a clearer value proposition, making it more compelling to first time shoppers. In contrast, segments like frequent buyers show only modest gains because the baseline messages are already well targeted; improvements here mainly come from tighter calls to action rather than drastic content changes. A few groups, for example price-drop seekers, record a slight dip in audience fit score, likely because the rewritten text trades some price sensitivity cues for broader appeal. Even so, their marketing scores still rise, indicating that the generated messages remain more persuasive overall despite minor targeting trade-offs.
To complement the evaluation of persuasive quality, we also measure how closely the generated messages preserve the meaning and style of the originals. Table 2 reports semantic and surface-level similarity scores using BERTScore-F1 (Zhang et al., 2019) and chrF (Popović, 2015), respectively. BERTScore-F1, computed with the multilingual xlm-roberta-large (Conneau et al., 2019) model, captures token-level semantic alignment between the original and generated messages. Meanwhile, chrF quantifies surface-level overlap based on character and word n-gram matches. The overall BERTScore-F1 of 0.87 and chrF of 24.22 suggest that while the rewrites are not mere copies, they still maintain strong semantic consistency with the source messages—allowing for effective reformulation without drifting from the original intent.
| Audience Segment | BERTScore-F1 | chrF |
| Potential New Customers | 0.87 | 23.44 |
| New Buyers | 0.87 | 23.15 |
| Lapsing Buyers | 0.87 | 21.24 |
| Abandoned Cart Buyers | 0.87 | 25.26 |
| Unpaid Order Buyers | 0.87 | 25.30 |
| Post-Purchase Group | 0.89 | 30.67 |
| Price-Drop Group | 0.89 | 26.13 |
| Active Old Followers | 0.87 | 20.48 |
| Frequent Buyers | 0.86 | 24.49 |
| New Followers | 0.88 | 22.80 |
| Repeat Buyers | 0.86 | 22.46 |
| Overall | 0.87 | 24.22 |
| Content | Template | Evaluate | Audience Score | Market Score | Preference | |||||
| Ori | Gen | (%) | Ori | Gen | (%) | Preferred | Rate (%) | |||
| GPT-4o | GPT-4o | GPT-4o | 4.18 | 4.56 | 9.09 | 3.33 | 4.61 | 38.44 | Gen | 78.44 |
| GPT-4o | GPT-4o + CoT | GPT-4o | 4.27 | 4.26 | -0.23 | 3.45 | 4.44 | 28.70 | Gen | 61.51 |
| GPT-4o | GPT-4o | GPT-4o + CoT | 3.80 | 4.30 | 13.16 | 3.25 | 4.49 | 38.15 | Gen | 81.17 |
| GPT-4o | GPT-4.1 mini | GPT-4o | 4.19 | 4.54 | 8.35 | 3.32 | 4.60 | 38.55 | Gen | 78.47 |
| GPT-4o | GPT-4o | GPT-4.1 mini | 4.20 | 4.43 | 5.48 | 3.30 | 4.61 | 39.70 | Gen | 81.05 |
| GPT-4o | GPT-4o | DeepSeek-R1 | 3.39 | 4.51 | 33.04 | 3.08 | 4.84 | 57.14 | Gen | 55.40 |
4.2 Ablation Studies
To assess how different LLM configurations affect system performance, we conduct an ablation study by varying the backbone models used in TemplateAgent and EvaluateAgent. Specifically, we explore combinations involving GPT-4o (OpenAI, 2024), GPT-4.1 mini (OpenAI, 2025a), and DeepSeek-R1 (Guo et al., 2025), as well as the use of chain-of-thought (CoT) (Wei et al., 2022) prompting. ContentAgent is kept fixed as GPT-4o to isolate the effects of template generation and evaluation components.
Table 3 shows that while the fully GPT-4o configuration delivers strong overall results, it is not uniformly superior. For example, applying chain-of-thought prompting to EvaluateAgent slightly boosts preference selection compared to the baseline, suggesting benefits in evaluation reasoning. In contrast, using CoT in TemplateAgent leads to degraded outcomes, likely due to verbose or unfocused generation. Models like GPT-4.1 mini offer performance comparable to GPT-4o, achieving similarly strong improvements in both audience and market scores, as well as high preference rates. This suggests that lighter-weight models can serve as viable drop-in replacements when resource constraints are a concern. In contrast, DeepSeek-R1 as EvaluateAgent yields the highest score improvements, but its relatively low preference rate indicates a potential mismatch between its scoring criteria and human-aligned quality judgments.
4.3 Case Studies
To illustrate how the CRMAgent adapts its messaging strategy across audience types, Table 4 presents two representative examples from potential new customers and abandoned cart buyers segments. In the first case, the original message is generic and celebratory, while the generated version introduces urgency (“Limited Time Offer”) and a clearer call to action, better aligning with the motivational needs of first-time shoppers. In the second example, the original offers a mild reminder, but the revised message adds a direct discount (“RM3 off”) and an upsell condition, using persuasive language to prompt checkout completion. These cases highlight how the agent enhances promotional effectiveness through more targeted and action-oriented rewrites.
| Case Study 1: Potential New Customers |
| Original: Everyone’s loving these! |
| Shoppers love these best-sellers—find your next favorite in our latest collection. |
| Generated: Limited Time Offer! |
| Don’t miss out on our hottest best-sellers! Explore our latest collection now and be the first to secure your favorites before they’re gone. Shop today! |
| Case Study 2: Abandoned Cart Buyers |
| Original: Did you forget something? |
| Looks like you left something in your cart. Ready to pick up where you left off? |
| Generated: Your cart is waiting—grab RM3 off now! |
| You left great picks in your cart! Complete your purchase with RM3 off when you spend RM15 or more. Don’t miss out—shop now and enjoy the benefits! |
5 Conclusion
In this paper, we introduce CRMAgent, a multi-agent system powered by LLMs for generating high-quality CRM message templates in the e-commerce domain. The system integrates three complementary strategies: content diagnosis, retrieval based adaptation, and rule based fallback to support merchants in crafting action oriented, persuasive copy across diverse audience segments. Extensive experiments show that CRMAgent consistently improves both audience alignment and marketing effectiveness compared to merchant written messages. In blind evaluations, the generated messages are preferred in the majority of cases, demonstrating the practical impact of our system. These findings suggest that multi-agent LLM systems can serve as scalable and effective solutions for automated CRM content creation, with potential to support merchants of varying expertise levels.
Limitations
While EvaluateAgent provides structured assessments of message quality, our evaluation is limited to offline settings. The generated templates have not yet been validated through online deployment or A/B testing, leaving real-world engagement impact unverified.
Appendix A Data
We begin with over 3 million successfully delivered CRM message logs from April 2025, progressively enriching each record with shop attributes, campaign metadata, product bindings, voucher information, and 7-day user engagement signals. This results in a detailed dataset indexed by plan, user, product, voucher, and template. We then group the data by key dimensions—including merchant, template, and voucher—to compute the average engagement score and sample count for each combination. In total, we obtain 15,806 aggregated records for downstream analysis.
Figure 7 illustrates the distribution of CRM plan volume across 11 predefined audience segments, each representing a distinct stage in the customer lifecycle. For example, “Potential New Customers” and “New Buyers” account for the largest shares, suggesting merchants often focus on early-stage acquisition. In contrast, groups such as “Repeat Buyers” and “Frequent Buyers” are comparatively smaller, indicating fewer efforts are directed toward re-engagement or loyalty building. This segmentation provides essential context for interpreting downstream engagement metrics, as user behavior and responsiveness tend to vary significantly across audience types. This breakdown helps ensure that engagement patterns are interpreted within the right audience context, rather than aggregating over inherently different customer types.
The engagement score behind Figure 8 reflects weighted user behaviors within seven days of receiving a message: one point for reading, three for clicking on a store or voucher, four for clicking a product card, two for interacting with CRM buttons, and minus five for unsubscribing. We compute this score for each unique combination of merchant, template, product, audience segment, and voucher setup. The resulting distribution is highly skewed—most plans yield low engagement, while a small number achieve high scores. Since engagement levels vary significantly across audience segments (e.g., abandoned-cart users tend to be more responsive than old followers), we analyze each segment separately. Within each group, we define the top 25% of campaigns as strong performers and the bottom 25% as weak ones—these serve as contrasting examples for downstream optimization.
Figure 9 shows the distribution of CRM plans across product categories. Many plans either lack a product binding (“None”) or focus on fast-moving goods like Beauty & Personal Care or Fashion Accessories. This category distribution influences not only engagement patterns but also determines how templates are generated. For instance, when no product is specified, the system falls back to rule-based generation during retrieval. In contrast, when product categories are available, we can use category-matching strategies. Thus, product type plays a central role in routing template generation.
Appendix B Agent Responses
In this section, we present illustrative examples to demonstrate how TemplateAgent and EvaluateAgent operate. The TemplateAgent generates improved versions of underperforming CRM messages by adjusting tone, structure, and content to better align with audience expectations. The EvaluateAgent then compares the original and rewritten messages, scoring each in terms of audience match and marketing effectiveness, and selecting the preferred one.
B.1 TemplateAgent Responses
To better understand how the TemplateAgent improves CRM messaging, we rewrite all message templates associated with poor-performing campaigns, as identified within each audience segment. Figures 10 and 11 illustrate representative examples of these rewrites for two specific audience types.
Figure 10 targets potential new customers who have not yet established a strong relationship with the brand and often require clear incentives to engage. The original message uses a soft, emotional tone (“we really missed you”) that, while friendly, lacks specificity and urgency. Such messaging may fail to capture the attention of users who are still undecided or unaware of the brand’s value proposition. In contrast, the rewritten version is more strategic: it introduces urgency with phrases like “limited-time steals” and emphasizes concrete benefits such as “guaranteed savings”. This not only strengthens the call to action but also aligns with the psychological triggers that are more effective for acquisition-focused messaging. The revised copy is concise, benefit-driven, and better tailored to convert first-time buyers by highlighting what they stand to gain immediately.
Figure 11 showcases a message tailored to active old followers, a segment that already has a relationship with the brand and tends to be more responsive to loyalty-oriented communication. The original message offers a vague incentive by mentioning “a special offer for your first order” and closes with a generic “shop now” prompt, which may feel impersonal or insufficient for users who have previously engaged. In contrast, the rewritten message opens with an explicit expression of appreciation and provides a clear benefit—twenty percent off the next order. This wording reinforces the value of continued engagement while making the reward tangible. The closing line, “before this offer disappears,” introduces urgency in a natural way, encouraging immediate action without sounding overly pushy. Overall, the revised message better aligns with the expectations and motivations of returning followers by combining gratitude, specificity, and timing.
B.2 EvaluateAgent Responses
To prevent agents from being biased by labels such as “original” or “generated”, each pair of CRM messages is anonymized as Message A and Message B during evaluation. While A and B correspond to the original and generated versions respectively, the evaluators are unaware of which is which—ensuring judgments focus solely on content quality.
For the potential new customer segment, Figure 12 illustrates a clear advantage for Message B in both audience relevance and marketing effectiveness. Message A assumes a prior relationship (“we really missed you”), which is inappropriate for new users. In contrast, Message B makes no such assumption and instead uses urgency (“limited availability”) and value (“guaranteed savings”) to attract first-time buyers. As a result, Message B receives the highest possible scores for both alignment and persuasion.
Similarly, in the active old followers segment shown in Figure 13, Message A misfires by offering a “first order” discount to users who are likely repeat customers. This mismatch undermines both credibility and clarity. Message B, on the other hand, acknowledges prior engagement and offers a concrete benefit (“20% off your next order”), paired with strong calls to action. It better suits the audience profile and earns top scores for both audience fit and marketing strength.
For the potential new customer segment, Figure 14 shows that EvaluateAgent strongly prefers Message B. This message avoids assuming any prior customer relationship and instead focuses on driving action through urgency (“limited-time steals”) and clarity (“guaranteed savings”). These are precisely the types of persuasive techniques that appeal to new customers who may be unfamiliar with the brand and require strong incentives to convert. In contrast, Message A’s tone—“we really missed you”—is clearly targeted at returning users and thus mismatched for this audience. The agent correctly identifies this discrepancy and selects the more appropriate and effective version.
In the case of active old followers (Figure 15), EvaluateAgent again prefers Message B for its audience fit and marketing clarity. Message B directly acknowledges the customer’s continued support and provides a well-specified incentive—“20% off your next order”—that feels both personal and actionable. Meanwhile, Message A refers vaguely to a “first order”, which can confuse or alienate users who are already loyal. The agent’s decision highlights the importance of aligning promotional language with customer lifecycle stage—what works for new users may backfire with returning ones.
Appendix C Template Quality Analyses
We conduct a systematic analysis of CRM message template quality using our custom evaluation sheet (Appendix D). Each template is assessed by o3 (OpenAI, 2025b) and assigned exactly one error type according to predefined criteria. Figure 16 presents the distribution of identified errors.
We find that the most common issue is Weak Call-to-Action (2,057 cases), where messages fail to provide a clear or urgent prompt for user action. Vague Incentive appears frequently as well (1,290 cases), often reflecting a lack of concrete or detailed value propositions. Misaligned Tone accounts for 570 cases, indicating mismatches in emotional or stylistic alignment between message and audience. Audience Assumption Errors and Irrelevant Offers occur less frequently but still signal problematic assumptions about customer history or behavior. This motivates the development of TemplateAgent, which is guided by these insights to generate templates that better match audience needs and communication goals.
Appendix D Instruction Sheet for Template Review
Provided Materials
Original Template: Each example includes only the original poor performance CRM message template.
Audience Segment: The target customer segment for the message is specified (e.g., potential new customers, active old followers).
Voucher: If applicable, the corresponding voucher information is provided (e.g., voucher title, discount type).
Product: If applicable, the corresponding promoted product information is included (e.g., product name or category).
Steps to Follow
Read the message carefully.
Categorize the problem using the error types below.
Provide short feedback explaining your choice.
Error Types and Criteria
Audience Assumption Error
Description: Assumes incorrect prior engagement with the user.
Indicators: References to past interaction for new customers; offers for first-time purchase to repeat customers.
Example: “Thanks for following!” sent to non-followers.
Weak Call-to-Action
Description: Lacks clarity or urgency in motivating the user to act.
Indicators: Generic phrases; no time-sensitive or action-oriented language.
Example: “Shop now” with no clear reason or benefit.
Vague Incentive
Description: Mentions a promotion without specifying what it is.
Indicators: Use of terms like “special offer” with no details on discount or benefit.
Example: “Here’s a deal for you!” without content.
Misaligned Tone
Description: The tone does not match the audience’s familiarity level.
Indicators: Overly casual or formal; assumes emotional connection without basis.
Example: “We miss you!” to someone who never interacted.
Irrelevant Offer
Description: The offer does not fit the customer’s likely status or behavior.
Indicators: Offering first-order discounts to frequent buyers.
Example: “Your first order!” for loyal customers.
Example Application
Audience Segment: Potential New Customer
Message: “It’s been a while—check out our bestsellers!”
Identified Error(s): Audience Assumption Error
Feedback: This implies a prior relationship which does not exist. Rewrite to welcome the new visitor and highlight value.
Audience Segment: Active Old Follower
Message: “Thanks for following—enjoy your first-order discount!”
Identified Error(s): Irrelevant Offer, Vague Incentive
Feedback: This audience is not new; avoid ‘first order’ wording. Clearly state what the benefit is.